log7 + log (n - 2) = log 6n
log516 - log52t = log52
log5m = log5125
logy = log16 + log49
log6(b2 + 2) + log62 = 2

logy = log16 + log49

Answer by mukhopadhyay(490)

log7 + log (n - 2) = log 6n
=> log [7(n-2) = log 6n
=> 7n-14 = 6n
=> n = 14;
.............
log5m = log5125
=> 5m = 5125
=> m = 1025
..................
logy = log16 + log49
=> y = (16)(49)
=> y = 784
...............
log6(b + 2) + log62 = 2
=> log(base6)[2(b+2)] = 2
=> 2(b+2) = 36
=> b=16
...............